OpenStudy (anonymous):

When solving the system by substitution, what would you plug into the second equation for the letter y? 2x+y=6 2x+2y=8 y = 2 y = 6-2x x = 6 x = 2y+6

OpenStudy (baldymcgee6):

2x+y=6 Solve for y by subtracting 2x from each side y = 6-2x

OpenStudy (asnaseer):

@baldymcgee6 is right. your last question used whats called the "elimination method" to solve these equations. here you are asked to use the "substitution method". in the substitution method, you first work out an expression for one of the variables from one of the equations, and then "substitute" this into the other equation to solve for the second variable.

OpenStudy (anonymous):

so y= 6-2x?

OpenStudy (anonymous):

cause thats what i guessed cause u subtract 2 from 2 and u get 0

OpenStudy (baldymcgee6):

That's correct

OpenStudy (anonymous):

so thats the answer (: yay!

OpenStudy (asnaseer):

not sure what you meant by "cause u subtract 2 from 2 and u get 0 "?

OpenStudy (anonymous):

well it said subtract 2 from 2

OpenStudy (anonymous):

2x from each side

OpenStudy (anonymous):

ohh i get it now!!!

OpenStudy (asnaseer):

good :)